# Privacy-Aware In-Context Learning for Large Language Models

### v3

@@ Abstract
Large language models (LLMs) have significantly transformed natural language understanding and generation, but they raise privacy concerns due to potential exposure of sensitive information. Studies have highlighted the risk of information leakage, where adversaries can extract sensitive information embedded in the prompts. In this work, we introduce a novel private prediction framework for generating high-quality synthetic text with strong privacy guarantees. Our approach leverages the Differential Privacy (DP) framework to ensure worst-case theoretical bounds on information leakage without requiring any fine-tuning of the underlying models. The proposed method performs inference on private records and aggregates the resulting per-token output distributions. This enables the generation of longer and coherent synthetic text while maintaining privacy guarantees. Additionally, we propose a simple blending operation that combines private and public inference to further enhance utility. Empirical evaluations demonstrate that our approach outperforms previous state-of-the-art methods on in-context-learning (ICL) tasks, making it a promising direction for privacy-preserving text generation while maintaining high utility.
@@ Introduction
However, this workflow has data privacy risks as sensitive records can be exposed in prompts and responses. For example, when LLMs are deployed with user data, such as clinical reports, incorporated into prompts, there is a risk that sensitive information can be inadvertently disclosed to non-relevant users [1]. An adversary can also extract sensitive user data in prompts via “jailbreaks”, where even entire prompts or segments of prompts can be extracted verbatim from some attacks. A mitigation approach could be to scrub Personally Identifiable Information (PII) from prompts, but even with that, there have been cases, such as in linkage attacks, where a combination of secondary information is enough to link the record back to an individual [2], leading to potential privacy violations.
Differential Privacy (DP) [3] has been previously used for protecting individual data as it ensures that sensitive information stays confidential while allowing meaningful insights to be drawn from the aggregated dataset. The U.S. Census Bureau’s LEHD OnTheMap tool [4], Google’s RAPPOR system as part of Google Chrome [5], Apple’s DP implementation [6, 7], Microsoft’s Telemetry collection [8], and Facebook and Social Science One’s release of election dataset [9, 10] are some noteworthy examples of industry-level deployment of this technology, whose reliability is reinforced by verification advances [11, 12]. To apply DP in LLM workflows, perhaps a simple and elegant way would be to transform the original prompt corpus into a semantically equivalent synthetic dataset. This synthetic version of the private dataset preserves the same overall patterns as the real data but contains no actual user records, making it safe to use for model training, inference, or external sharing without risking privacy breaches.
Current approaches to creating differentially private text with large language models fall into two main groups: private fine-tuning and private prediction [13]. Fine-tuning methods update model weights on the private data using a DP-SGD style algorithm [14]. Once fine-tuned, the model generates synthetic text directly. This method often produces high-quality outputs but requires extensive compute for training as well as full access to the model parameters. On the other hand, private prediction based methods only rely on test-time inference instead of fine-tuning the model [15, 16]. In these approaches, noise is added to the model’s output distribution so that each generated token satisfies differential privacy. As they avoid any form of model training, synthetic examples can be produced on demand. However, because a separate privacy cost is incurred for each generated token, the overall privacy budget accumulates rapidly which limits this approach to generating a smaller amount of synthetic text [17].
@@
We focus on the privacy-preserving in-context learning (ICL) framework, where large language models (LLMs) can perform downstream tasks effectively using only a few demonstrations, without requiring fine-tuning, as demonstrated in [18]. Among the notable efforts in privacy-preserving ICL, [19] introduce a differentially private (DP) inference mechanism by constructing a consensus over ensembles of queries with disjoint demonstrations. Although their method satisfies DP guarantees, it incurs a privacy cost for each query, thereby limiting the number of queries that can be answered under a given privacy budget. [20] also explore private ensembling but rely on the availability of unlabeled public data, which is labeled using a teacher ensemble via ICL. This reliance on public data contrasts with our approach, which operates solely on private data, making it more suitable for sensitive domains such as healthcare or industrial applications where public datasets with similar distributions may not be accessible. Furthermore, both works primarily target text classification, while our method extends to a broader range of tasks, as demonstrated in our experiments.
@@
Our contribution aligns with the general domain of synthetic text generation under privacy constraints, diverging from approaches that rely heavily on private fine-tuning [21, 22]. Inspired by the capabilities of LLMs [18], our method leverages their generation abilities in a DP-compliant manner by privately aggregating generation probabilities over disjoint subsets of private demonstrations. This technique draws conceptual similarity to the PATE framework [23, 24], which generates private models by training on public data labeled by an ensemble of teacher models trained on disjoint private data. Extensions of PATE to text generation include SeqPATE [25] and Submix [26], which introduce domain-specific adaptations.
@@
Another direction is DP decoding, as proposed by [16], which combines LLM predictions with uniform distributions. However, these methods generally require private training on the sensitive data, unlike our lightweight approach. Compared to private fine-tuning techniques for synthetic data generation [27, 28, 29, 30], our method avoids computationally intensive fine-tuning and is well-suited for scenarios where few-shot ICL suffices. While private fine-tuning may be preferable for generating large volumes of synthetic data, our approach balances efficiency and effectiveness in low-data settings.
@@
Beyond ICL, privacy-preserving efforts in natural language generation encompass several techniques. Word-level noise injection and metric local differential privacy (LDP) have been used for sanitized text generation [31, 32, 33, 34]. Other methods, such as those by [35] and [36], apply LDP directly to full documents through fine-tuning or zero-shot prompting followed by sanitization. Private fine-tuning remains prominent in synthetic data generation: [37] utilize DP-SGD [14] to fine-tune LLMs, while [38] demonstrate improvements via parameter-efficient fine-tuning like LoRA [39]. Two-stage fine-tuning approaches have also been proposed [40], and similar ideas have been extended to structured data [41]. Another line of work focuses on private prediction [13], where privacy is guaranteed only for outputs, often via subsample-and-aggregate techniques [42], as used in PATE [43]. Applied to synthetic text, these ideas involve per-token privacy accounting [17, 44], but suffer from limited utility due to the high privacy cost of each token. Other adaptations of private prediction to LLMs [45, 26, 46] have not focused on synthetic generation. Lastly, private filtering methods operate on entire LLM responses and rely on matching public data via embedding similarity or keyword selection [47, 48, 49], but lack adaptability to new data distributions.
@@
In-context learning (ICL) leverages a pre-trained model to utilize its existing knowledge by conditioning it on a sequence of demonstration examples without any further gradient updates to the model weights [50, 51]. During inference, the model receives several such input–label demonstration pairs which follow a consistent format, followed by a novel test input subjected to the same pattern. The model has to then autoregressively predict the correct label for the final prompt in a few-shot manner, effectively learning the task from the provided context rather than solely relying on the information stored in model parameters [52]. Recently, ICL has proven to be versatile across diverse NLP tasks ranging from text classification and question answering, especially as model scale increases, highlighting the emergent capabilities of LLMs [53].
@@
Definition 1 (Differential Privacy (DP) [54]).
@@ 9 $x_{t}\sim\operatorname{softmax}(\hat{z}/\tau)$
We present a private prediction protocol for next-token prediction. Our approach follows a two-step framework for in-context-learning (ICL) with differential privacy:
@@
Below, we begin by outlining standard LLM inference in a typical decoder based model, then introduce our differentially private prediction method for generating synthetic examples. Finally, we present the formal privacy guarantees provided by PRISM.
@@
Any decoder-only LLM such as GPT [55] and Llama [56] takes an input prompt and generates a sequence of token indices in auto-regressive manner. The model maps each generated token $x_{t}$ to a logit vector $z\in\mathbb{R}^{V}$, where $V$ is the token vocabulary size. This process involves initializing the prompt sequence $X$ with the instruction phrase $p$, and repeating the following steps: (a) compute logits for the next token $x_{t}$ as $z_{t}=\text{LLM}(x_{t})$, (b) sample the next token $x_{t}\sim\operatorname{softmax}(z_{t}/\tau)$ for the temperature hyper-parameter $\tau>0$, and (c) append $x_{t}$ to $X$. The process stops when $x_{t}$ is the end-of-sequence token <eos>, indicating the end of the response. Here, $\operatorname{softmax}(z_{t}/\tau)$ is the distribution that assigns probability proportional to $\exp(z_{t}/\tau)$ to the $t_{th}$ token, and $\tau>0$ is a hyper-parameter that flattens or sharpens the distribution.
@@
[17] 4 GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$ -
[57] 4 Gemma 1.1 2B IT $\mathbf{76.8_{4.8}}$ $72.3_{2.5}$ $38.8_{6.0}$ $47.7_{2.5}$ $81.7_{2.4}$
@@
11 [17] 4 GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
[17] 4 Gemma 1.1 2B IT $74.9_{3.8}$ $\mathbf{80.9_{3.6}}$ $36.7_{2.2}$ $34.1_{9.3}$ $78.7_{1.9}$
[57] 4 Gemma 1.1 2B IT $75.9_{3.5}$ $75.1_{0.5}$ $39.2_{3.7}$ $47.1_{6.0}$ $\mathbf{84.5_{1.0}}$
@@
TABLE I: In-context-learning accuracy comparison where we report mean and standard deviation over three random samplings (equally many from each label for classification; fully random for extraction) of synthetic/real data. (*) Note: For the results using GPT-3 babbage only the top-100 logprobs for contextual calibration (only top 5 are available now) are used. While not directly comparable to Gemma model which uses logprobs over the full vocabulary, we report their results for context similar to [57]. Best results for $\epsilon=\infty,\text{and }1$ on Gemma 1.1 2B IT are in bold.
@@
Since the averaged logit vector is generated from private subset hence each token selected from this vector adds to the privacy budget. To minimize the privacy leakage, we also use generate an axillary token distribution from the same LLM without access to the sensitive data. It uses a public prompt function that generates text for a given text of category $y_{i}$. At each iteration, a public token $u$ is generated by combining its clipped logits with the aggregated private logits $\bar{z}$ using simple averaging. According to our privacy analysis, clipping ensures that the influence of any individual private prompt is bounded by $\frac{c}{s}$ in each coordinate. clipping ensures that the influence of any individual private prompt is bounded by, this contribution is further reduced to $\frac{c}{2s}$. This reduction in sensitivity allows us to inject less noise and operate under a smaller privacy budget compared to [57]. From a utility perspective, since both $\bar{z}$ and $u$ are clipped in a rank‐preserving manner, their arithmetic mean preserves token preferences common to both sources. As a result, tokens strongly supported by both private and public contexts receive the highest scores, while those favored by only one are suppressed.
Lemma 1 (Exponential Mechanism [58]).
@@
At each iteration, we ensure differential privacy by selecting the new token using the exponential mechanism. Furthermore, using the composition property of DP [3], we guarantee that the entire sequence of upto $T$ generated tokens remains collectively is also differentially private.
@@
In this privacy bound, $c$ denotes the clipping threshold, $T$ is the number of composition steps, $s$ is the size of the sensitive subset, and $\tau$ is the temperature parameter used in softmax sampling. We provide the proof of Theorem 1 in the Appendix, which utilizes the utility guarantee of the exponential mechanism [58] and advanced composition for the exponential mechanism [3].
Experimental Setup
Datasets:
For evaluation, we follow the prior ICL work [59] and use the following setup. For classification tasks, we use three datasets: the 4-way news classification dataset AGNews [60], the 6-way question classification dataset TREC [60], and the 14-way topic classification DBPedia [61]. We also evaluate on two information extraction tasks, namely MIT-G, and MIT-D [62]. These are both slot-filling datasets with movie genre (MIT-G) and director name (MIT-D) as the slots to be filled. An illustrative example is shown in Figure 2.
Since our method requires access to the full token probability distribution at each decoding step, we use the instruction-tuned (IT) variant of Gemma 1.1 [63], a decoder-only LLM with two billion parameters, aligning with prior work by [57]. However, this choice is made purely for experimental benchmarking. Our algorithm is model-agnostic and can be applied to any decoder-style LLM, provided access to full token-level logit distribution is available. We provide additional details for chosen hyper-parameters in the Appendix.
Results
Table I presents our in‐context learning results on the five benchmark tasks. Our primary choice of LLM is Gemma-1.12B IT which provides access to model logits as opposed to closed models such as “GPT3-baggage” used by some previous works [17]. We evaluate and compare the utility of our model with varying level of privacy budget ($\varepsilon=0,\infty,1$). The first block with $\varepsilon=0$ has the highest amount of privacy shows zero-shot performance (no ICL) is uniformly poor accross tasks (e.g., 24.8% on AGNews).
The second block with $\varepsilon=\infty$ examines when the least amount of noise is added in our DP framework. We observe that we perform better than the current SOTA [57]’s baseline on Gemma-1.12B in four tasks, namely DBPedia, TREC, MIT-G and MIT-D, while lagging behind by only $3.3\%$ on AGNews. For most datasets: DBPedia, TREC, MIT-G and MIT-D, we also outperform [17]’s baseline with GPT3 Babbage that uses only top-100 logprobs.
The third block with $\varepsilon=1$ imposes stronger privacy constraints on synthetic data. We observe that we outperform both baselines by [17], and  [57] with Gemma-1.12B on three out of five tasks: AGNews, TREC, MIT-G. Here also, for AGNews, TREC, MIT-G and MIT-D datasets, we outperform [17]’s baseline with GPT3 Babbage that uses only top-100 logprobs.
Effects of K-Shot and Privacy Budget Settings
$\epsilon$ 1-shot 2-shot 4-shot 8-shot 12-shot
1 81.66 82.46 76.80 74.82 74.46
4 79.90 82.62 79.84 76.04 68.52
8 77.00 83.34 78.62 74.90 68.56
$\infty$ 80.86 80.48 80.94 81.44 75.48
TABLE II: ICL accuracy under varying privacy budgets with different K-shots on DBPedia.
$\epsilon$ 1-shot 2-shot 4-shot 8-shot 12-shot
1 68.32 60.60 63.00 64.40 65.84
4 68.76 59.72 62.48 63.08 63.28
8 68.96 62.00 63.20 67.08 62.32
$\infty$ 62.40 54.04 61.92 67.16 70.12
TABLE III: ICL accuracy under varying privacy budgets with different K-shots on TREC.
In Table II, results on DBPedia show that performance generally improves with larger values of $\epsilon$, indicating that relaxing privacy constraints allows for more effective use of private data in generating synthetic demonstrations. However, as the number of shots increases, the performance can sometimes degrade slightly possible due to the lack of targeted private signal which can also be the training artifacts of LLM training. Finally, the best performance for 4-shot (80.94), 8-shot (81.44), and 12-shot (75.48) appears in the $\epsilon=\infty$ (non-private) case, highlighting the utility-privacy trade-off inherent to DP algorithms.
Table III presents the corresponding analysis on the TREC dataset, where the pattern across privacy levels is more nuanced. For example, $\epsilon=8$ outperforms all other settings for it should be 1-shot, 2-shot, and 4-shot configurations, indicating that a moderately relaxed privacy guarantee yields significant utility benefits in certain tasks.
Structured Data Generation
We evaluate how well our approach preserves privacy while generating syntactic structures from sensitive data. In structured generation tasks, many tokens are essential for maintaining the correct output structure. To assess this, we experiment on the WikiMoviesJSON generation task, using preprocessing and evaluation setups described in [57]. We assess performance using two metrics: (1) the percentage of outputs that are syntactically well-formed JSON parses, and (2) the percentage of outputs that pass basic schema validation. As shown in Table IV, our method achieves high-quality and schema-compliant JSON generation even under a strict privacy budget of $\epsilon=1$, demonstrating the effectiveness of our approach for privacy-preserving structured text synthesis.
$\varepsilon$ Method $\tau$ Parses (%) Validates (%) #\#raw
1 [57] 2 $80.6_{1.3}$ $74.2_{1.9}$ $94.3_{1.2}$
2.5 $4.9_{1.1}$ $1.5_{0.1}$ $138.0_{7.5}$
Ours 1.13 $84.2_{4.08}$ $81.1_{8.1}$ $10.3_{1.15}$
TABLE IV: Results for generating JSON records from WikiMoviesJSON and report gains in structure preservation and validations. We report mean and standard deviation over 3 runs of dataset generation. Here, $\tau$ refers to the sampling temperature, and #\#raw refers to the number of raw samples produced before parsing and validation checks.
Privacy Attacks
While differential privacy offers theoretical privacy guarantees, empirical valid is essential [64]. We conduct personally identifiable information (PII) extraction attack on the Enron email dataset following the experimental setup described in [65]. Specifically, we first construct a private dataset of text which contain email addresses in the body. Using Algorithm 1, we we construct a private subset $S$ and generate $T=15$ synthetic tokens using the prompt template as “Extract only the email address from the above text.”. The goal of this attacks is to test whether our approach prevents release of private email addresses in the generated text. We evaluate under various privacy budget namely $\epsilon=[1,4,8]$. Across all tested privacy levels, we observe zerozero sensitive email address in the generated responses demonstrating the effectiveness of our approach in preventing leakage of PII and demonstrate its strong practical privacy-preserving capabilities.
Furthermore, to assess the practical privacy of DP few-shot generation for in-context learning, we conduct membership inference attacks (MIA) following [20]. When actual private samples are used in prompts, attacks succeed with high AUC (94.20 for $\epsilon=\infty$). In contrast, our DP approach significantly reduces the AUC ($51.78$ for $\epsilon=4$), confirming improved membership privacy. See Appendix Table MIA for details.
Discussion
Traditional differentially private generation methods, such as those employed by [57], rely on metric-based mechanisms like the Sparse Vector Technique (SVT), which requires threshold computations under additive noise. These approaches typically employ distributional distance metrics such as $\ell_{1}$ to score or select candidate tokens. However, these metrics are computed over normalized probability distributions (i.e., post-softmax), where even semantically similar tokens with slightly different probabilities or indices can yield high $\ell_{1}$ distance values. This issues is especially pronounced in high-dimensional output spaces of language models, where such metrics treat tokens as orthogonal dimensions, ignore semantic similarity, and are insensitive to token ranking that are crucial for meaningful text generation.
Moreover, approach by [57] continues generating additional sentences beyond termination, often yielding low-quality synthetic examples. Our procedure strictly terminates when <eos> token is encountered. This guarantees that all outputs are semantically coherent and suitable for downstream tasks. Qualitative examples supporting this claim are provided in the Appendix.
Conclusion
As access to foundation models grows, the resources required for training these models have become expensive; hence, private prediction could emerge as a compelling alternative to private fine-tuning. In this work, we show that private prediction can generate synthetic text while following the standard differential-privacy guarantees. This privately generated corpus substantially boosts performance in many-shot in-context learning. Moreover, introducing a mechanism for sampling tokens from public models and blending them with private tokens ehances the utility of prediction tasks without compromising the privacy accounting.
Limitations
Acknowledgment
Appendix A Proof of Theorem 1
@@
In this we compute the sensitivity of several functions used in Algorithm 1. Each function depends on a set of logit vectors. Recall that a logit vector $z$ is an element of $\mathbb{R}^{v}$. For any function $l$ defined as:
@@
where $\bar{z}=\frac{1}{s}\sum_{z\in Z}\text{clip}_{c}(z)$ is obtained by aggregating private logits and $u$ is a public logit vector.
@@
$\leq\frac{c}{2s}$
Composition
Lemma 3 (Advanced Composition for Exponential Mechanism  [3]).

### v4

# Privacy Preserving In-Context-Learning Framework for Large Language Models

@@ Abstract
Large language models (LLMs) have significantly transformed natural language understanding and generation, but they raise privacy concerns due to potential exposure of sensitive information. Studies have highlighted the risk of information leakage, where adversaries can extract sensitive information embedded in the prompts. In this work, we introduce a novel private prediction framework for generating high-quality synthetic text with strong privacy guarantees. Our approach leverages the Differential Privacy (DP) framework to ensure worst-case theoretical bounds on information leakage without requiring any fine-tuning of the underlying models. The proposed method performs inference on private records and aggregates the resulting per-token output distributions. This enables the generation of longer and coherent synthetic text while maintaining privacy guarantees. Additionally, we propose a simple blending operation that combines private and public inference to further enhance utility. Empirical evaluations demonstrate that our approach outperforms previous state-of-the-art methods on in-context-learning (ICL) tasks, making it a promising direction for privacy-preserving text generation while maintaining high utility. Our code is available at https://github.com/bhusalb/privacy-preserving-icl.
@@ Introduction
However, this workflow has data privacy risks as sensitive records can be exposed in prompts and responses. For example, when LLMs are deployed with user data, such as clinical reports, incorporated into prompts, there is a risk that sensitive information can be inadvertently disclosed to non-relevant users (wang2023decodingtrust). An adversary can also extract sensitive user data in prompts via “jailbreaks”, where even entire prompts or segments of prompts can be extracted verbatim from some attacks. A mitigation approach could be to scrub Personally Identifiable Information (PII) from prompts, but even with that, there have been cases, such as in linkage attacks, where a combination of secondary information is enough to link the record back to an individual (powar2023sok), leading to potential privacy violations.
Differential Privacy (DP) (dwork2014algorithmic) has been previously used for protecting individual data as it ensures that sensitive information stays confidential while allowing meaningful insights to be drawn from the aggregated dataset. The U.S. Census Bureau’s LEHD OnTheMap tool (census), Google’s RAPPOR system as part of Google Chrome (rappor), and Facebook and Social Science One’s release of election dataset (DVN/TDOAPG_2020; Evans_King_2023) are some noteworthy examples of industry-level deployment of this technology. To apply DP in LLM workflows, perhaps a simple and elegant way would be to transform the original prompt corpus into a semantically equivalent synthetic dataset. This synthetic version of the private dataset preserves the same overall patterns as the real data but contains no actual user records, making it safe to use for model training, inference, or external sharing without risking privacy breaches.
Current approaches to creating differentially private text with large language models fall into two main groups: private fine-tuning and private prediction (dwork2018privacy). Fine-tuning methods update model weights on the private data using a DP-SGD style algorithm (abadi2016deep). Once fine-tuned, the model generates synthetic text directly. This method often produces high-quality outputs but requires extensive compute for training as well as full access to the model parameters. On the other hand, private prediction based methods only rely on test-time inference instead of fine-tuning the model (van2020trade; majmudar2022differentially). In these approaches, noise is added to the model’s output distribution so that each generated token satisfies differential privacy. As they avoid any form of model training, synthetic examples can be produced on demand. However, because a separate privacy cost is incurred for each generated token, the overall privacy budget accumulates rapidly which limits this approach to generating a smaller amount of synthetic text (tang2024privacypreserving).
@@
We focus on the privacy-preserving in-context learning (ICL) framework, where large language models (LLMs) can perform downstream tasks effectively using only a few demonstrations, without requiring fine-tuning, as demonstrated in BrownMRSK20. Among the notable efforts in privacy-preserving ICL, panda2023differentially introduce a differentially private (DP) inference mechanism by constructing a consensus over ensembles of queries with disjoint demonstrations. Although their method satisfies DP guarantees, it incurs a privacy cost for each query, thereby limiting the number of queries that can be answered under a given privacy budget. duan2023flocks also explore private ensembling but rely on the availability of unlabeled public data, which is labeled using a teacher ensemble via ICL. This reliance on public data contrasts with our approach, which operates solely on private data, making it more suitable for sensitive domains such as healthcare or industrial applications where public datasets with similar distributions may not be accessible. Furthermore, both works primarily target text classification, while our method extends to a broader range of tasks, as demonstrated in our experiments.
@@
Our contribution aligns with the general domain of synthetic text generation under privacy constraints, diverging from approaches that rely heavily on private fine-tuning (yu2022differentially; LiTLH21). Inspired by the capabilities of LLMs (BrownMRSK20), our method leverages their generation abilities in a DP-compliant manner by privately aggregating generation probabilities over disjoint subsets of private demonstrations. This technique draws conceptual similarity to the PATE framework (PapernotAEGT17; PapernotSMRTE18), which generates private models by training on public data labeled by an ensemble of teacher models trained on disjoint private data. Extensions of PATE to text generation include SeqPATE (tian2022seqpate) and Submix (ginart2022submix), which introduce domain-specific adaptations.
@@
Another direction is DP decoding, as proposed by majmudar2022differentially, which combines LLM predictions with uniform distributions. However, these methods generally require private training on the sensitive data, unlike our lightweight approach. Compared to private fine-tuning techniques for synthetic data generation (yue2023synthetic; mattern-etal-2022-differentially; mireshghallah2023; carranza2023privacypreserving), our method avoids computationally intensive fine-tuning and is well-suited for scenarios where few-shot ICL suffices. While private fine-tuning may be preferable for generating large volumes of synthetic data, our approach balances efficiency and effectiveness in low-data settings.
@@
Beyond ICL, privacy-preserving efforts in natural language generation encompass several techniques. Word-level noise injection and metric local differential privacy (LDP) have been used for sanitized text generation (feyisetan2020privacy; xu-etal-2020-differentially; carvalho2021tem; du2023sanitizing). Other methods, such as those by mattern-etal-2022-limits and utpala2023locally, apply LDP directly to full documents through fine-tuning or zero-shot prompting followed by sanitization. Private fine-tuning remains prominent in synthetic data generation: yue2022synthetic utilize DP-SGD (abadi2016deep) to fine-tune LLMs, while kurakin2023harnessing demonstrate improvements via parameter-efficient fine-tuning like LoRA (hu2021lora). Two-stage fine-tuning approaches have also been proposed (wu2024prompt), and similar ideas have been extended to structured data (tran2024differentially). Another line of work focuses on private prediction (dwork2018privacy), where privacy is guaranteed only for outputs, often via subsample-and-aggregate techniques (nissim2007smooth), as used in PATE (papernot2018scalable). Applied to synthetic text, these ideas involve per-token privacy accounting (tang2024privacypreserving; hong2023dp), but suffer from limited utility due to the high privacy cost of each token. Other adaptations of private prediction to LLMs (majmudar2022differentially; ginart2022submix; flemings2024differentially) have not focused on synthetic generation. Lastly, private filtering methods operate on entire LLM responses and rely on matching public data via embedding similarity or keyword selection (yu2024privacy; xie2024differentially; wu2023privacy), but lack adaptability to new data distributions.
@@
In-context learning (ICL) leverages a pre-trained model to utilize its existing knowledge by conditioning it on a sequence of demonstration examples without any further gradient updates to the model weights (brown2020language; liu2021pretrainpromptpredictsystematic). During inference, the model receives several such input–label demonstration pairs which follow a consistent format, followed by a novel test input subjected to the same pattern. The model has to then autoregressively predict the correct label for the final prompt in a few-shot manner, effectively learning the task from the provided context rather than solely relying on the information stored in model parameters (dong2024surveyincontextlearning). Recently, ICL has proven to be versatile across diverse NLP tasks ranging from text classification and question answering, especially as model scale increases, highlighting the emergent capabilities of LLMs (wei2022emergent).
@@
Definition 1 (Differential Privacy (DP) (DworkKMMN06)).
@@ 9 $x_{t}\sim\operatorname{softmax}(\hat{z}/\tau)$
We present a private prediction protocol for next-token prediction in the Figure 1. Our approach follows a two-step framework for in-context-learning (ICL) with differential privacy:
@@
Below, we begin by outlining standard LLM inference in a typical decoder based model, then introduce our differentially private prediction method for generating synthetic examples. Finally, we present the formal privacy guarantees provided by our algorithm.
@@
Any decoder-only LLM such as GPT (gpt) and Llama (liu2022llama) takes an input prompt and generates a sequence of token indices in auto-regressive manner. The model maps each generated token $x_{t}$ to a logit vector $z\in\mathbb{R}^{V}$, where $V$ is the token vocabulary size. This process involves initializing the prompt sequence $X$ with the instruction phrase $p$, and repeating the following steps: (a) compute logits for the next token $x_{t}$ as $z_{t}=\text{LLM}(x_{t})$, (b) sample the next token $x_{t}\sim\operatorname{softmax}(z_{t}/\tau)$ for the temperature hyper-parameter $\tau>0$, and (c) append $x_{t}$ to $X$. The process stops when $x_{t}$ is the end-of-sequence token <eos>, indicating the end of the response. Here, $\operatorname{softmax}(z_{t}/\tau)$ is the distribution that assigns probability proportional to $\exp(z_{t}/\tau)$ to the $t_{th}$ token, and $\tau>0$ is a hyper-parameter that flattens or sharpens the distribution.
@@
tang2024privacypreserving 4 GPT-3 babbage $69.3_{4.8}$ $82.3_{3.7}$ $50.6_{6.9}$ $54.4_{7.0}$ -
amin2024private 4 Gemma 1.1 2B IT $\mathbf{76.8_{4.8}}$ $72.3_{2.5}$ $38.8_{6.0}$ $47.7_{2.5}$ $81.7_{2.4}$
@@
11 tang2024privacypreserving 4 GPT-3 babbage $64.1_{3.9}$ $81.2_{3.0}$ $50.7_{4.1}$ $46.3_{7.8}$ $69.2_{7.9}$
tang2024privacypreserving 4 Gemma 1.1 2B IT $74.9_{3.8}$ $\mathbf{80.9_{3.6}}$ $36.7_{2.2}$ $34.1_{9.3}$ $78.7_{1.9}$
amin2024private 4 Gemma 1.1 2B IT $75.9_{3.5}$ $75.1_{0.5}$ $39.2_{3.7}$ $47.1_{6.0}$ $\mathbf{84.5_{1.0}}$
@@
Table 1: In-context-learning accuracy comparison where we report mean and standard deviation over three random samplings (equally many from each label for classification; fully random for extraction) of synthetic/real data. (*) Note: For the results using GPT-3 babbage only the top-100 logprobs for contextual calibration (only top 5 are available now) are used. While not directly comparable to Gemma model which uses logprobs over the full vocabulary, we report their results for context similar to (amin2024private). Best results for $\epsilon=\infty,\text{and }1$ on Gemma 1.1 2B IT are in bold.
@@
Since the averaged logit vector is generated from private subset hence each token selected from this vector adds to the privacy budget. To minimize the privacy leakage, we also use generate an axillary token distribution from the same LLM without access to the sensitive data. It uses a public prompt function that generates text for a given text of category $y_{i}$. At each iteration, a public token $u$ is generated by combining its clipped logits with the aggregated private logits $\bar{z}$ using simple averaging. According to our privacy analysis, clipping ensures that the influence of any individual private prompt is bounded by $\frac{c}{s}$ in each coordinate and after merging with public logits this is further reduced to $\frac{c}{2s}$. This reduction in sensitivity allows us to inject less noise and operate under a smaller privacy budget compared to amin2024private. From a utility perspective, since both $\bar{z}$ and $u$ are clipped in a rank‐preserving manner, their arithmetic mean preserves token preferences common to both sources. As a result, tokens strongly supported by both private and public contexts receive the highest scores, while those favored by only one are suppressed.
Lemma 1 (Exponential Mechanism (mcsherry2007mechanism)).
@@
At each iteration, we ensure differential privacy by selecting the new token using the exponential mechanism. Furthermore, using the composition property of DP (dwork2014algorithmic), we guarantee that the entire sequence of upto $T$ generated tokens remains collectively is also differentially private.
@@
In this privacy bound, $c$ denotes the clipping threshold, $T$ is the number of composition steps, $s$ is the size of the sensitive subset, and $\tau$ is the temperature parameter used in softmax sampling.
@@
We analyze the sensitivity of several functions used in Algorithm 1. Each such function is defined on a multiset $Z=\{z_{1},z_{2},\dots,z_{n}\}\subset\mathbb{R}^{v}$, where each $z_{i}\in\mathbb{R}^{v}$ is a logit vector. Consider a the function $\ell:(\mathbb{R}^{v})^{n}\to\mathbb{R}^{v}$ corresponding to line 8 of Algorithm 1 defined as:
@@
where $u\in\mathbb{R}^{v}$ is a public logit vector, and $\bar{z}=\frac{1}{s}\sum_{z\in Z}\text{clip}_{c}(z)$ is the mean of the clipped private logits. The operator $\text{clip}_{c}(\cdot)$ performs element-wise clipping with threshold $c>0$, and $s$ is the number of logit vectors in the private logit vectors.
@@
$\leq\frac{c}{2s}\qed$
Lemma 3 (Advanced Composition for Exponential Mechanism  (dwork2014algorithmic)).